OpenStudy (anonymous):

The surface area of a sphere is found using the formula , where r is the radius. If the surface area of a sphere is 64 square feet, what is the radius, in feet?

OpenStudy (anonymous):

surface area of sphere \[ A=4\pi r ^{2}\] 64=4 times pi times r r=64/4times pi r=64/4(3.142) r=5.092 feet please dont forget to give this as best response